# Tensor network influence functionals in the continuous-time limit: connections to quantum embedding, bath discretization, and higher-order time propagation

### v1

@@ III.1 Boundary Influence functional tensor
Defining the continuous-time limit of the influence functional tensor network is in principle one way to eliminate the time-step error from the standard second-order Trotter decomposition, and in numerical applications allows for the introduction of a wide variety of higher-order differential integrators. However, as shown in [33], the IF-MPS shows a nonphysical entanglement entropy scaling in the limit of $\Delta t\to 0$, as the entanglement entropy always scales to zero. This suggests that the continuous-time limit requires a more careful treatment.
@@
Consider the continuous-time limit $\Delta t\to 0$, where $\hat{U}^{SB}$ can be expressed as,
@@
and the corresponding matrix elements of the IF-MPS, $A^{s_{i},s_{f}}$, are,
$\begin{gathered}A^{0,0}=I-i\hat{L}_{B}\Delta t\\ A^{0,1}=A^{1,0}=-i\hat{L}_{SB}\Delta t\\ A^{1,1}=I-i\hat{L}_{B}\Delta t.\end{gathered}$ (32)
These do not have the same form as the tensor entries in a cMPS [53, 54], which take the following form,
$\begin{gathered}A^{0}=I+\epsilon Q\\ A^{1}=\sqrt{\epsilon}R\\ A^{2}=\frac{1}{2}\epsilon R^{2},\end{gathered}$ (33)
@@
There are two main differences between Eq.  32 and Eq.  33. First, the IF has an additional ‘$I$’ term in $A^{1,1}$, that is not in $A^{2}$ in Eq. 33. Second, the $A^{1}$ terms are proportional to $\Delta t$, not $\sqrt{\Delta t}$. It is clear to see the effect of these two differences in the formulation of the differential equation of motion for the 1-RDM $\Gamma^{R}$. In the infinitesimal limit of Eq. 26, after expanding $U^{SB}$ the first term becomes
@@
This corresponds to unitary dynamics in the super-Fock bath orbital space, which preserves the spectrum of $\Gamma^{R}$. The initial $\Gamma^{R}$ is given by the pure state $\mathinner{|{\rho_{B}}\rangle\!\rangle}$, so its single-particle spectrum consists of only 0 and 1. Thus, the entanglement entropy of the IF is zero. Note that this result is the manifestation of the fact that the term in Eq. 35 is proportional to $(\Delta t)^{2}$, instead of $(\Delta t)^{1}$, which implies that the bath dynamics does not have effective ‘dissipation’ terms.
@@
$\begin{gathered}t=USV=(US^{1/2})(S^{1/2}V)=t^{S}t^{B}\\ \hat{O}_{a}^{S}=\sum_{s}t^{S}_{sa}\hat{d}_{s}^{{\dagger}},\quad\hat{O}_{a}^{B}=\sum_{i}t^{B}_{ai}\hat{a}_{i},\end{gathered}$ (38)
where $n_{S}$ is the number of singular values. This leads us to write $\hat{U}^{SB}$ as follows,
@@
$\begin{gathered}\hat{W}^{S}_{a}=\begin{cases}\hat{I}_{S}&a=0\\ \hat{O}^{S}_{a}\sqrt{\Delta t}&1\leq a\leq n_{S}\\ \hat{O}^{S{\dagger}}_{a-n_{S}}\sqrt{\Delta t}&n_{S}+1\leq a\leq 2n_{S}\end{cases}\\ \hat{W}^{B}_{a}=\begin{cases}\hat{I}_{B}-i\hat{L}_{B}\Delta t&a=0\\ -i\hat{O}^{B}_{a}\sqrt{\Delta t}&1\leq a\leq n_{S}\\ -i\hat{O}^{B{\dagger}}_{a-n_{S}}\sqrt{\Delta t}&n_{S}+1\leq a\leq 2n_{S}\end{cases}\end{gathered}$ (40)
Figure 5: The time evolution operator, $\hat{U}^{SB}$ can be split into two tensors, $\hat{W}^{S}$ and $\hat{W}^{B}$, which can be expressed as $\hat{U}^{SB}=\sum_{a}\hat{W}^{S}_{a}\otimes\hat{W}^{B}_{a}$. In the single impurity case with two impurity orbitals in the Liouville space, the index $a$ has four components.
@@
The matrix elements of the boundary IF-MPS are then given by,
$\begin{gathered}W^{B,0}={I}_{B}-i{L}_{B}\Delta t\\ W^{B,1}=-i\sqrt{\Delta t}\sum_{a}\left(f_{a}^{{\dagger}}O_{a}^{B}+\text{h.c.}\right)\\ W^{B,2}=\Delta t\sum_{a}|\hat{O}_{a}^{B}|^{2},\end{gathered}$ (42)
which satisfies the form of Eq. 33  [^1] The factor of 1/2 difference from $W^{B,2}$ comes from the fact that Eq. 33 describes double excitations of indistinguishable bosonic particles whereas $W^{B,2}$ takes into account different auxiliary fermions.. Therefore the boundary IF-MPS has a well-defined continuum limit. It is also possible to construct $\hat{W}^{B}$ from $\hat{U}^{SB}$ for general $\Delta t$ outside the continuous-time limit, which is described in the Supplemental Material. Henceforth, we will implicitly assume that the IF-MPS refers to the boundary IF-MPS except where the distinction is important.
@@ III.2 Gauge dynamics in the continuous-time IF-MPS
The boundary IF-MPS matrix elements define a new effective time evolution operator in the Liouville space of the system and bath with a well-defined continuous-time limit with non-vanishing entanglement entropy. This makes it possible to express and compress the MPS in the continuous-time limit.
As described in Sec. II.4, in the case of a non-interacting bath, the object that defines the conversion to canonical form and the compression is the 1-RDM of the right/left IF state, $\Gamma^{R}$ and $\Gamma^{L}$. Using $\hat{W}^{B}$, we can construct the corresponding iterative procedure to update the 1-RDM of the right/left IF state. In the continuous-time limit, $\Delta t\rightarrow 0$, this leads to an equation of motion for the 1-RDM of the right/left IF state.
The equation of motion for the 1-RDM of the right IF state, $\Gamma^{R}$, is given by,
@@
where $\{A,B\}=AB+BA$ is an anti-commutator. A detailed derivation of this equation of motion is in the Supplemental Material. It clearly shows that $t^{B}$ acts as a dissipation term. Similarly, the equation of motion for the 1-RDM of the left IF state, $\Gamma^{L}$, is given by,
@@
From $\Gamma^{L}$, we can find the gauge transformation $\hat{G}$. This is then absorbed into the right IF state to canonicalize the IF-MPS. For ${\Gamma}^{G}$, the 1-RDM of the gauge-transformed right IF state, we first define the gauge-transformed time evolution operator
@@
where we set the $\dot{R}^{L}$ terms to be zero when $k=l$ and $|\nu_{k}-\nu_{l}|<\varepsilon$. Note that $R^{L}$ is only defined up to degenerate eigenvectors of $\Gamma^{L}$, therefore, to fix this redundancy, we propagate $R^{L}$ with the regularized $\dot{R}^{L}$ from a reference time, which we set to be the final time of the $\Gamma^{L}$ propagation.
With this gauge-transformed time evolution operator, the equation of motion for ${\Gamma}^{G}$ can be written as follows.
@@
After obtaining $\Gamma^{G}$ in this continuous-time picture, the effective bath orbitals and their rotation matrix, $R^{\text{eff}}$ (see Sec. II.4), can be defined from the eigenvectors of $\Gamma^{G}$. $R^{\text{eff}}$ is then used to construct the time-dependent projection operator, $\mathcal{P}$, which defines the compression of the IF-MPS in the continuous-time setting.
@@ III.3 Embedding Liouville operator for the embedded wavefunction
Once we solve the equation of motion for ${\Gamma}^{G}(t)$, its truncated spectrum defines a set of effective bath orbitals at all (continuous) times $t$. By projecting the original Liouville dynamics into the time-dependent embedding space (of system and bath orbitals) and retaining terms first order in $\Delta t$, we can extract the generator for the embedded wavefunction, which we call here the embedding Liouville operator, $\hat{L}^{\text{emb}}$. This yields a continuous-time state-vector propagation governed by the boundary IF-MPS.
First, by expanding Eq. 30, we have $\mathcal{P}\hat{{L}}_{G}\mathcal{P}$ where $\mathcal{P}$ is the projection operator into the effective bath orbitals and $\hat{{L}}_{G}=\hat{G}\hat{L}_{SB}\hat{G}^{-1}+\hat{{L}}_{GB}$. In addition, there is also a term for the time dependence of the effective bath orbitals in the projection operator. Using the rotation matrix of the effective bath orbitals, $R^{\text{eff}}$, the additional Liouville operator from this time dependence can be written as,
@@
where $m$ and $n$ index the effective bath orbital basis. Finally, the impurity-only Liouville operator can be simply added. The embedding Liouville operator becomes,
@@
We can now write down the equation of motion for the embedded wavefunction, $\mathinner{|{\Psi^{\text{emb}}}\rangle\!\rangle}$, with which we can carry out state-vector propagation,
@@
Note that the embedding Liouville operator, $\hat{L}^{\text{emb}}$, is time-dependent.
@@ III.4 Connections to embedding theories
where $\hat{X}$ is a quadratic Hermitian operator that describes the time dependence of the effective bath orbitals, which has the same form as Eq. 48, but here, its elements are determined from the TDVP equations. The equation of motion in Eq. 52 defines the embedding Hamiltonian $\hat{H}+\hat{X}$ which corresponds to the embedding Liouville operator, $\hat{L}^{\text{emb}}$, in Eq. 49. Aside from the different generators of the bath dynamics, a qualitative difference between real-time DMET and the embedding scheme that derives from the IF-MPS in this work is the nonunitary nature of the gauge transformation, $\hat{G}$. In the real-time DMET, the $\hat{X}$ operator implements unitary dynamics of the effective bath orbitals.
@@ IV Results
Figure 6: The time-dependence of the double occupancy of the impurity, $\langle n_{\uparrow}n_{\downarrow}\rangle$, in the quench dynamics of the symmetric Anderson model with $U=2.5\pi\Gamma$ and $\varepsilon_{\sigma}=-1.25\pi\Gamma$. The impurity is quenched from an initial unoccupied state and a thermal bath of temperature $\Gamma\beta=2$. Reference results from time-dependent DMRG are shown (tdDMRG, black dotted). (a) Results from the discrete-time boundary IF-MPS using different numbers of effective bath orbitals, $N_{\text{eff}}=$ 4 (cyan), 6 (green), 8 (blue), and 10 (red) with a time-step of $\Gamma\Delta t=0.01$. This shows that at $N_{\text{eff}}=8$ and $10$, the IF-MPS dynamics are fully converged to the eye compared to the reference results. (b) Results from the time-independent Lanczos-based bath discretization with thermofield transformation using a series of static bath discretizations, $N_{b}^{\text{th}}=$ 4 (cyan), 6 (green), 8 (blue), and 10 (red), where $N_{b}^{\text{th}}$ is the number of Lanczos vectors in the thermofield transformed bath. The finite size bath errors are clearly larger than in (a).
We now describe simulations that implement the above boundary IF-MPS formulation, including its continuous-time formulation, in the single impurity Anderson model. To generate reference results, we used state-vector propagation with a large bath discretization of the AIM. Specifically, we used $N_{b}=40$ discrete bath orbitals ($N_{b}$ refers to orbitals of each spin, i.e. 40 spin $\uparrow$ and 40 spin $\downarrow$ orbitals) to approximate the bath spectral density function, $J(\omega)$, using a linear scheme [59, 3], as follows
@@
where $W=10\Gamma$, $\omega\in[-W,W]$, and $I_{i}=[-W+\frac{2W}{N_{b}}(i-1),-W+\frac{2W}{N_{b}}i]$, and the impurity Hamiltonian parameters were chosen as $U=-2\varepsilon_{\sigma}=2.5\pi\Gamma$. The initial bath state was taken to be the (decoupled) thermal state with $\Gamma\beta=2$, and the impurity was quenched from an initial unoccupied state, $\hat{\rho}_{S}(0)=\mathinner{|{0}\rangle}\!\!\mathinner{\langle{0}|}$. Benchmark results were then computed using the time-dependent density matrix renormalization group (tdDMRG) with bond dimension up to 600 using the Block2 [60, 61] package. We carried out a thermofield transformation on the bath orbitals to reduce the bond dimension of the MPS, as described in [5]. We then propagated the MPS using the two-site time-dependent variational principle. With the largest bond dimension 600 and a time-step size of $\Gamma\Delta t=0.01$, all the local observables presented here are converged with respect to bond dimension and time-step to within an estimated absolute error of $10^{-5}$.
@@
For the continuous-time implementation, we wrote the effective $\hat{L}^{\text{emb}}$ as a second-quantized fermionic operator which could then be used in state-vector propagation using a higher-order integrator such as 4th-order Runge-Kutta. Note that to apply 4th-order Runge-Kutta to the state-vector at time $t$, with time-step $\Delta t$, requires $\hat{L}^{\text{emb}}(t)$ at the intermediate time $t+\Delta t/2$. To obtain this, we propagated the bath 1-RDM equations of motion (Eqs. 4344) with a finer time-step $\Delta t_{\text{fine}}$, also with the 4th-order Runge-Kutta integrator. Since the cost to propagate the bath 1-RDM is much lower than that for the many-body wavefunction, we used $\Delta t_{\text{fine}}$ much smaller than $\Delta t$ to minimize the time-step error from the bath 1-RDM propagation. We fixed $\Gamma\Delta t_{\text{fine}}=0.01/2^{6}\approx 0.00016$ in this work. We then used the quantum chemistry full configuration interaction (exact diagonalization) implementation in PySCF [64, 65] to carry out the propagation of the state-vector.
One assumption for higher-order numerical propagators to be accurate is that the time-dependent embedding Liouville operator, $\hat{L}^{\text{emb}}$, is well-behaved, i.e. it does not change too sharply. However, in the initial and final periods of the time propagation, we observe that the spectrum of the Liouville operator diverges. This is because $\Gamma^{L}$ consists of nearly core and virtual orbitals at either temporal boundary and large values of $g_{k}$ and $g_{k}^{-1}$ in the gauge transformation are applied to the effective bath orbital spaces.
This issue can easily be circumvented by initially propagating the wavefunction without the gauge transformation for a short time. In this short time period, the effective bath orbitals are taken from $\Gamma^{R}$ and $\Gamma^{L}$. After the completion of this short time period, the gauge transformation is applied to the wavefunction, and subsequently, the wavefunction is propagated with the gauge-transformed equation of motion. We found that the spectrum of the Liouville operator showed instabilities up to an initial $\Gamma\Delta t_{i}=0.06$ and after a final $\Gamma\Delta t_{f}=0.1$ (Supplemental Material). We therefore used propagation without gauges in these short periods, and between these boundary times, the wavefunction was propagated with the gauge transformation.

### v2

@@ III.1 Boundary influence functional tensor
Defining the continuous-time limit of the influence functional tensor network is in principle one way to eliminate the time step error from the standard second-order Trotter decomposition, and in numerical applications allows for the introduction of a wide variety of higher-order differential integrators. However, as shown in [37], the IF-MPS shows a nonphysical entanglement entropy scaling in the limit of $\Delta t\to 0$, as the entanglement entropy always scales to zero. This suggests that the continuous-time limit requires a more careful treatment.
@@
Consider the continuous-time limit $\Delta t\to 0$, where $\hat{U}^{SB}$ can be expressed as
@@
and the corresponding matrix elements of the IF-MPS $A^{s_{i},s_{f}}$ are,
$\begin{gathered}A^{0,0}=I-i\hat{L}_{B}\Delta t,\\ A^{0,1}=A^{1,0}=-i\hat{L}_{SB}\Delta t,\\ A^{1,1}=I-i\hat{L}_{B}\Delta t.\end{gathered}$ (32)
These do not have the same form as the tensor entries in a cMPS [59, 60], which take the following form:
$\begin{gathered}A^{0}=I+\epsilon Q,\\ A^{1}=\sqrt{\epsilon}R,\\ A^{2}=\frac{1}{2}\epsilon R^{2},\end{gathered}$ (33)
@@
There are two main differences between Eq.  32 and Eq.  33. First, the IF has an additional “$I$” term in $A^{1,1}$, that is not in $A^{2}$ in Eq. 33. Second, the $A^{1}$ terms are proportional to $\Delta t$, not $\sqrt{\Delta t}$. It is clear to see the effect of these two differences in the formulation of the differential equation of motion for the 1-RDM $\Gamma^{R}$. In the infinitesimal limit of Eq. 26, after expanding $U^{SB}$ the first term becomes
@@
This corresponds to unitary dynamics in the super-Fock bath orbital space, which preserves the spectrum of $\Gamma^{R}$. The initial $\Gamma^{R}$ is given by the pure state $\mathinner{|{\rho_{B}}\rangle\!\rangle}$, so its single-particle spectrum consists of only 0 and 1. Thus, the entanglement entropy of the IF is zero. Note that this result is the manifestation of the fact that the term in Eq. 35 is proportional to $(\Delta t)^{2}$, instead of $(\Delta t)^{1}$, which implies that the bath dynamics does not have effective “dissipation” terms.
@@
$\begin{gathered}t=USV=(US^{1/2})(S^{1/2}V)=t^{S}t^{B},\\ \hat{O}_{a}^{S}=\sum_{s}t^{S}_{sa}\hat{d}_{s}^{{\dagger}},\quad\hat{O}_{a}^{B}=\sum_{i}t^{B}_{ai}\hat{a}_{i},\end{gathered}$ (38)
where $n_{S}$ is the number of singular values. This leads us to write $\hat{U}^{SB}$ as follows:
@@
$\begin{gathered}\hat{W}^{S}_{a}=\begin{cases}\hat{I}_{S}&a=0\\ \hat{O}^{S}_{a}\sqrt{\Delta t}&1\leq a\leq n_{S}\\ \hat{O}^{S{\dagger}}_{a-n_{S}}\sqrt{\Delta t}&n_{S}+1\leq a\leq 2n_{S},\end{cases}\\ \hat{W}^{B}_{a}=\begin{cases}\hat{I}_{B}-i\hat{L}_{B}\Delta t&a=0\\ -i\hat{O}^{B}_{a}\sqrt{\Delta t}&1\leq a\leq n_{S}\\ -i\hat{O}^{B{\dagger}}_{a-n_{S}}\sqrt{\Delta t}&n_{S}+1\leq a\leq 2n_{S}.\end{cases}\end{gathered}$ (40)
Figure 5: The time-evolution operator, $\hat{U}^{SB}$ can be split into two tensors, $\hat{W}^{S}$ and $\hat{W}^{B}$, which can be expressed as $\hat{U}^{SB}=\sum_{a}\hat{W}^{S}_{a}\otimes\hat{W}^{B}_{a}$. In the single impurity case with two impurity orbitals in the Liouville space, the index $a$ has four components.
@@
The matrix elements of the boundary IF-MPS are then given by
$\begin{gathered}W^{B,0}={I}_{B}-i{L}_{B}\Delta t,\\ W^{B,1}=-i\sqrt{\Delta t}\sum_{a}\left(f_{a}^{{\dagger}}O_{a}^{B}+\text{H.c.}\right),\\ W^{B,2}=\Delta t\sum_{a}|\hat{O}_{a}^{B}|^{2},\end{gathered}$ (42)
which satisfies the form of Eq. 33  [^2] The factor of 1/2 difference from $W^{B,2}$ comes from the fact that Eq. 33 describes double excitations of indistinguishable bosonic particles whereas $W^{B,2}$ takes into account different auxiliary fermions.. Therefore, the boundary IF-MPS has a well-defined continuum limit. It is also possible to construct $\hat{W}^{B}$ from $\hat{U}^{SB}$ for general $\Delta t$ outside the continuous-time limit, which is described in the SM. Henceforth, we will implicitly assume that the IF-MPS refers to the boundary IF-MPS except where the distinction is important.
@@ III.2 Gauge dynamics in the continuous-time IF-MPS
The boundary IF-MPS matrix elements define a new effective time-evolution operator in the Liouville space of the system and bath with a well-defined continuous-time limit with nonvanishing entanglement entropy. This makes it possible to express and compress the MPS in the continuous-time limit.
As described in Sec. II.4, in the case of a noninteracting bath, the object that defines the conversion to canonical form and the compression is the 1-RDM of the right/left IF state, $\Gamma^{R}$ and $\Gamma^{L}$. Using $\hat{W}^{B}$, we can construct the corresponding iterative procedure to update the 1-RDM of the right/left IF state. In the continuous-time limit, $\Delta t\rightarrow 0$, this leads to an equation of motion for the 1-RDM of the right/left IF state.
The equation of motion for the 1-RDM of the right IF state $\Gamma^{R}$ is given by
@@
where $\{A,B\}=AB+BA$ is an anticommutator. A detailed derivation of this equation of motion is in the SM. It clearly shows that $t^{B}$ acts as a dissipation term. Similarly, the equation of motion for the 1-RDM of the left IF state $\Gamma^{L}$ is given by
@@
From $\Gamma^{L}$, we can find the gauge transformation $\hat{G}$. This is then absorbed into the right IF state to canonicalize the IF-MPS. For ${\Gamma}^{G}$, the 1-RDM of the gauge-transformed right IF state, we first define the gauge-transformed time-evolution operator
@@
where we set the $\dot{R}^{L}$ terms to be zero when $k=l$ and $|\nu_{k}-\nu_{l}|<\varepsilon$. Note that $R^{L}$ is only defined up to degenerate eigenvectors of $\Gamma^{L}$; therefore, to fix this redundancy, we propagate $R^{L}$ with the regularized $\dot{R}^{L}$ from a reference time, which we set to be the final time of the $\Gamma^{L}$ propagation.
With this gauge-transformed time-evolution operator, the equation of motion for ${\Gamma}^{G}$ can be written as follows:
@@
After obtaining $\Gamma^{G}$ in this continuous-time picture, the effective bath orbitals and their rotation matrix $R^{\text{eff}}$ (see Sec. II.4) can be defined from the eigenvectors of $\Gamma^{G}$. $R^{\text{eff}}$ is then used to construct the time-dependent projection operator $\mathcal{P}$, which defines the compression of the IF-MPS in the continuous-time setting.
@@ III.3 Embedding Liouville operator for the embedded wavefunction
Once we solve the equation of motion for ${\Gamma}^{G}(t)$, its truncated spectrum defines a set of effective bath orbitals at all (continuous) times $t$. By projecting the original Liouville dynamics into the time-dependent embedding space (of system and bath orbitals) and retaining terms first order in $\Delta t$, we can extract the generator for the embedded wavefunction, which we call here the embedding Liouville operator $\hat{L}^{\text{emb}}$. This yields a continuous-time state vector propagation governed by the boundary IF-MPS.
First, by expanding Eq. 30, we have $\mathcal{P}\hat{{L}}_{G}\mathcal{P}$ where $\mathcal{P}$ is the projection operator into the effective bath orbitals and $\hat{{L}}_{G}=\hat{G}\hat{L}_{SB}\hat{G}^{-1}+\hat{{L}}_{GB}$. In addition, there is also a term for the time dependence of the effective bath orbitals in the projection operator. Using the rotation matrix of the effective bath orbitals $R^{\text{eff}}$, the additional Liouville operator from this time dependence can be written as
@@
where $m$ and $n$ index the effective bath orbital basis. Finally, the impurity-only Liouville operator can be simply added. The embedding Liouville operator becomes
@@
We can now write down the equation of motion for the embedded wavefunction $\mathinner{|{\Psi^{\text{emb}}}\rangle\!\rangle}$ with which we can carry out state vector propagation,
@@
Note that the embedding Liouville operator $\hat{L}^{\text{emb}}$ is time-dependent.
@@ III.4 Connections to embedding theories
where $\hat{X}$ is a quadratic Hermitian operator that describes the time dependence of the effective bath orbitals, which has the same form as Eq. 48, but here, its elements are determined from the TDVP equations. The equation of motion in Eq. 52 defines the embedding Hamiltonian $\hat{H}+\hat{X}$ which corresponds to the embedding Liouville operator $\hat{L}^{\text{emb}}$ in Eq. 49. Aside from the different generators of the bath dynamics, a qualitative difference between real-time DMET and the embedding scheme that derives from the IF-MPS in this paper is the nonunitary nature of the gauge transformation $\hat{G}$. In the real-time DMET, the $\hat{X}$ operator implements unitary dynamics of the effective bath orbitals.
@@ IV Results
Figure 6: The time-dependence of the double occupancy of the impurity $\langle n_{\uparrow}n_{\downarrow}\rangle$ in the quench dynamics of the symmetric Anderson model with $U=2.5\pi\gamma$ and $\varepsilon_{\sigma}=-1.25\pi\gamma$. The impurity is quenched from an initial unoccupied state and a thermal bath with two different temperatures, $\gamma\beta=2$ (a, b) and $\gamma\beta=50$ (c, d). Reference results from time-dependent DMRG are shown (tdDMRG, black dotted). (a, c) Results from the discrete-time boundary IF-MPS using different numbers of effective bath orbitals, $N_{\text{eff}}=$ 4 (cyan), 6 (green), 8 (blue), and 10 (red) with a time step of $\gamma\Delta t=0.01$. (b, d) Results from the time-independent Lanczos-based bath discretization with thermofield transformation using a series of static bath discretizations, $N_{b}^{\text{th}}=$ 4 (cyan), 6 (green), 8 (blue), and 10 (red), where $N_{b}^{\text{th}}$ is the number of Lanczos vectors in the thermofield transformed bath. The finite size bath errors are clearly larger than in (a) and (c).
We now describe simulations that implement the above boundary IF-MPS formulation, including its continuous-time formulation, in the single impurity Anderson model. To generate reference results, we used state vector propagation with a large bath discretization of the AIM. Specifically, we used $N_{b}=40$ discrete bath orbitals ($N_{b}$ refers to orbitals of each spin, i.e., 40 spin $\uparrow$ and 40 spin $\downarrow$ orbitals) to approximate the bath spectral density, $J(\omega)=\sum_{i}|t_{i}|^{2}\delta(\omega-E_{i})$, using a linear scheme [65, 7], as follows
@@
where $W=10\gamma$, $\omega\in[-W,W]$, and $I_{i}=[-W+\frac{2W}{N_{b}}(i-1),-W+\frac{2W}{N_{b}}i]$, and the impurity Hamiltonian parameters were chosen as $U=-2\varepsilon_{\sigma}=2.5\pi\gamma$. The initial bath state was taken to be the (decoupled) thermal state with two different temperature regimes: 1) An intermediate-temperature regime with $\gamma\beta=2$ where the temperature is comparable to the other energy scales, and 2) a low-temperature regime with $\gamma\beta=50$, which is a lower temperature than the Kondo temperature, estimated to be $T_{K}\sim 0.07\gamma$ [9] ($\Gamma\beta_{K}=\gamma/T_{K}\sim 14$). The impurity was quenched from an initial unoccupied state, $\hat{\rho}_{S}(0)=\mathinner{|{0}\rangle}\!\!\mathinner{\langle{0}|}$. Benchmark results were then computed using the time-dependent density matrix renormalization group (tdDMRG) with bond dimensions up to 300 using the Block2 [66, 67] package. We carried out a thermofield transformation on the bath orbitals to reduce the bond dimension of the MPS, as described in [9]. We then propagated the MPS using the two-site time-dependent variational principle. With the largest bond dimension 300 and a time step size of $\gamma\Delta t=0.01$, all the local observables presented here are converged with respect to bond dimension and time step to within an estimated absolute error of $10^{-5}$.
@@
For the continuous-time implementation, we wrote the effective $\hat{L}^{\text{emb}}$ as a second-quantized fermionic operator which could then be used in state vector propagation using a higher-order integrator such as 4th-order Runge-Kutta. Note that to apply 4th-order Runge-Kutta to the state vector at time $t$, with time step $\Delta t$, requires $\hat{L}^{\text{emb}}(t)$ at the intermediate time $t+\Delta t/2$. To obtain this, we propagated the bath 1-RDM equations of motion (Eqs. 4344) with a finer time step $\Delta t_{\text{fine}}$, also with the 4th-order Runge-Kutta integrator. Since the cost to propagate the bath 1-RDM is much lower than that for the many-body wavefunction, we used $\Delta t_{\text{fine}}$ much smaller than $\Delta t$ to minimize the time step error from the bath 1-RDM propagation. We fixed $\gamma\Delta t_{\text{fine}}=0.01/2^{6}\approx 0.00016$ in this paper. We then used the quantum chemistry full configuration interaction (exact diagonalization) implementation in PySCF [70, 71] to carry out the propagation of the state vector.
One assumption for higher-order numerical propagators to be accurate is that the time-dependent embedding Liouville operator $\hat{L}^{\text{emb}}$ is well-behaved, i.e., it does not change too sharply. However, in the initial and final periods of the time propagation, we observe that the spectrum of the Liouville operator diverges. This is because $\Gamma^{L}$ consists of nearly core and virtual orbitals at either temporal boundary and large values of $g_{k}$ and $g_{k}^{-1}$ in the gauge transformation are applied to the effective bath orbital spaces.
This issue can easily be circumvented by initially propagating the wavefunction without the gauge transformation for a short time. In this short time period, the effective bath orbitals are taken from $\Gamma^{R}$ and $\Gamma^{L}$. After the completion of this short time period, the gauge transformation is applied to the wavefunction, and subsequently, the wavefunction is propagated with the gauge-transformed equation of motion. We found that the spectrum of the Liouville operator showed instabilities up to an initial $\gamma\Delta t_{i}=0.06$ and after a final $\gamma\Delta t_{f}=0.1$ (see the SM). We therefore used propagation without gauges in these short periods, and between these boundary times, the wavefunction was propagated with the gauge transformation.